21. In one-factor ANOVA for independent samples, which of the following reflects a random variability of scores in the samples?

  • A. MSbetween

  • B. MStotal

  • C. MSwithin

  • D. A & B

22. A one-factor ANOVA is used to evaluate the mean differences in cognitive task performance between three experimental groups, with n = 15 participants in each group. If the computed F = 3.17, which of the following is the correct statistical decision?

  • A. Reject the null hypothesis with either α = .05 or α = .01.

  • B. Reject the null hypothesis with α = .05 but fail to reject with α = .01.

  • C. Fail to reject the null hypothesis with either α = .05 or α = .01.

  • D. Fail to reject the null hypothesis with α = .05 but reject with α = .01.

24.A researcher uses a one-factor ANOVA to evaluate the mean difference between two treatments and obtains F(1, 22) = 4.00. If the researcher used a t-test for independent samples to evaluate the data, what would be the computed t statistic?

  • A. 2.00

  • B. 8.00

  • C. 16.00

  • D. Not enough information given to answer this question.

25. In a one-factor ANOVA for independent samples, the larger the differences between sample means,______________________________.

  • A. the smaller the numerator of F-ratio will be.

  • B. the smaller the numerator of F-ratio will be.

  • C. the larger the numerator of F-ratio will be.

  • D. the larger the denominator of F-ratio will be.
Answer #1

21)

C. MSwithin

22)

df numerator =number of groups k-1=3-1=2

df denominator=nk-k=3*15-3 =42

p value from excel =fdist(3.17,2,42) =0.0522

C. Fail to reject the null hypothesis with either α = .05 or α = .01.

24)

since t test statistic =√F =√4 =2

A. 2.00

25)

C. the larger the numerator of F-ratio will be.